OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 15 15 over 2 9 11 Please explain, I am not understanding this.

OpenStudy (anonymous):

Look : if we have : 2*(2+3)=2*2+2*3 or : 2(a+8)=2*a+2*8=2a+16 k ? So : − 4(y + 1) = (-4)*y+(-4)*+1=-4y-4 ( look : (-)*(-)=+ (+)*(+)=+ (_)*(+)=- ) so if we have : 4 (2x − 14)=4*(2x)+4*(-14)=8x-56 and we had : 1/4 (2x − 14) = 1/8x-56=4 if we have a Multiplication ( * 8x-56) 8x-56*(1/8x-56)= 8x-56*(+4)=>1=32x-56*(+4)=>1=32x-224 225=32x x=225/32= Got it ?;)

OpenStudy (anonymous):

i think my way was right but idk why my answer isnt there ! Please check your question again ... but i think this way can help you in another questions ;)